A station supplies 250KVA at a lagging power factor of 0.8. A synchronious motor is connected in parallel with the load. If the combined load is 250KW with a lagging P.F of 0.9, determine: 

1. The leading KVAR taken by the motor

2. The KVA rating of the motor

3. P.F at which the motor operates


1
Expert's answer
2021-05-31T06:20:49-0400

Load = - 250 kVA,

0.8 lagging power "\phi" = 36.86'

PL = (250 x 0.8) kW = 200 kW

QL = (250 x sin 36.86) = 150 kVAR Combined Load = - 250 kW,

0.9 lagging "\phi" = 25.84;

QTotal = P x tan "\phi" = (250 x tan 25.84 ) kVAR = 121 kVAR

QTotal = (QLoad + Qmotor)

121 = 150 + (-Qmotor)

Qmotor = 29 kVAR

PTotal = (PLoad + Pmotor)

250 = 200 + Pmotor

Pmotor = 50 kW

s = (P2 + Q2)1/2 kVA = (502 x 292)1/2

s = 57.80 kVA

cos "\phi"m = P/s = 50/57.80

cos "\phi"m = 0.865